if i am going 75mph and going 1000 feet, how many seconds would i travel 1000 feet going 75mph
Answer by Alan3354(30983)

if i am going 75mph and going 1000 feet, how many seconds would i travel 1000 feet going 75mph
------------
60 mi/hr = 88 ft/sec (I remember that, you can work it out if you like.)
75 = 60*1.25
75 mi/hr = 88*1.25 ft/sec = 110 ft/sec
-----
1000/110 =~ 9.091 seconds